# DocMath-Eval: Evaluating Math Reasoning Capabilities of LLMs in Understanding Long and Specialized Documents

### v2

@@ Abstract
Recent LLMs have demonstrated remarkable performance in solving exam-like math word problems. However, the degree to which these numerical reasoning skills are effective in real-world scenarios, particularly in expert domains, is still largely unexplored. This paper introduces DocMath-Eval, a comprehensive benchmark specifically designed to evaluate the numerical reasoning capabilities of LLMs in the context of understanding and analyzing specialized documents containing both text and tables. We evaluate a wide spectrum of 48 LLMs with Chain-of-Thought and Program-of-Thought prompting methods, aiming to comprehensively assess the capabilities and limitations of existing LLMs in DocMath-Eval. We found that even the current best-performing system (i.e., GPT-4o) still significantly lags behind human experts in solving complex numerical reasoning problems grounded in long contexts. We believe that DocMath-Eval can serve as a valuable benchmark for evaluating LLMs’ capabilities in solving challenging numerical reasoning problems within expert domains.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have attracted significant attention due to their capabilities in solving a broad range of tasks OpenAI (2022, 2023); Touvron et al. (2023), including math word problems (MWPs) commonly found in academic exams Wang et al. (2017); Miao et al. (2020); Amini et al. (2019); Cobbe et al. (2021); Hendrycks et al. (2021); Cobbe et al. (2021); Lu et al. (2023); Chen et al. (2023b). These MWPs vary from basic arithmetic to advanced algebra, showcasing LLMs’ proficiency in numerical reasoning — a crucial skill for interpreting and manipulating numerical data across various contexts. Despite this progress, there is still a significant gap in understanding the practicality of LLMs’ numerical reasoning in real-world scenarios, particularly in specialized fields such as finance, medicine, and science. As illustrated in Figure 1, these expert domains necessitate LLMs to interpret complex, domain-specific documents, applying numerical reasoning to complex problem-solving Chen et al. (2021); Zhu et al. (2021); Zhao et al. (2022); Li et al. (2022b). Recognizing this gap, our research focuses on the finance domain Li et al. (2022a); Wu et al. (2023a); Yang et al. (2023); Callanan et al. (2023); Xie et al. (2024). The finance industry often deals with lengthy and data-intensive documents that demand advanced numerical reasoning skills for accurate analysis and decision-making.
@@
We evaluate a wide spectrum of open- and closed-source LLMs, specifically, 48 models from 17 organizations. Two prompting methods, Chain-of-Thought (CoT) Wei et al. (2022) and Program-of-Thought (PoT) Chen et al. (2023a), are adopted for experiments. Our experimental results indicate that while the existing best-performing LLM on average (i.e., GPT-4o) can achieve high performance in simple settings (e.g., DM), it still falls short of human experts in more challenging ones, i.e.,, DM. Moreover, Claude-3.5-Sonnet outperforms other LLMs, achieving an accuracy of 40.0% on the DM set when applying CoT prompting. However, it still lags far behind human expert performance, which stands at 76%. This significant gap between LLMs and human experts underscores the challenges presented by DocMath-Eval. It highlights the need for further advancements in adapting LLMs’ numerical reasoning capabilities for practical application in real-world expert domains.
@@ 2 Related Work
Numerical reasoning over documents requires models to have a deep understanding of context and the ability to derive answers through numerical reasoning Dua et al. (2019). Applying these models in the finance domain Xie et al. (2023); Wu et al. (2023a); Yang et al. (2023) presents additional challenges in terms of interpreting hybrid data Zhu et al. (2021) and utilizing domain-specific expertise Chen et al. (2021); Zhao et al. (2023a). Numerous datasets focusing on numerical reasoning over specialized documents have been proposed recently. Two notable benchmarks are TAT-QA Zhu et al. (2021) and FinQA Chen et al. (2021), which represent pioneering efforts in studying numerical reasoning in finance, particularly requiring the fusion of tabular and textual content. Building upon TAT-QA, a more challenging dataset named TAT-HQA Li et al. (2022b) was developed, focusing on counterfactual questions in relation to the provided context. Additionally, MultiHiertt Zhao et al. (2022) focuses on numerical reasoning over longer financial documents containing multiple tables. However, as illustrated in Table 1, these four datasets focus on less challenging scenarios, where either simple numerical reasoning (e.g., calculating the increasing rate or average value) is sufficient, or the input context is short. Furthermore, there is a lack of a standardized benchmark for systematically evaluating models’ performance across varying difficulty levels in terms of numerical reasoning and document understanding.
@@ 3.2 Solution Format Standardization
We observe that existing finance QA datasets feature solutions in various formats. Specifically, TAT-QA Zhu et al. (2021) and TAT-HQA Li et al. (2022b) utilize text, while MultiHiertt Zhao et al. (2022) employs mathematical expressions, such as 100/3, and FinQA Chen et al. (2021) uses math programs, such as divide(100,3), for solution annotations. This diversity in annotation formats hinders the development of a unified evaluation framework to assess LLM performance across different benchmarks. Additionally, solutions in text format often lack the precision and unambiguity necessary for computational problem-solving; and solutions in mathematical equations or math programs are less descriptive, with the semantic meaning of the equations sometimes being unclear.
To overcome the aforementioned limitations, in DocMath-Eval, we represent solutions using Python programs. Such a unified Python program format supports a standardized and effective evaluation framework for LLM assessment. Specifically, annotators are required to first define variables at the beginning of the Python function, starting with “def solution():”. These variables correspond to the key elements or quantities mentioned in the question or question-relevant content in the documents. Annotators are instructed to assign meaningful names that clearly represent each element. They then write a sequence of Python statements that logically solve the problem, step by step. Additionally, annotators receive a bonus for writing detailed comments, thereby enhancing the code’s readability and understandability. To ensure the accuracy and functionality of the solutions, our annotation interface automatically executes the Python function. This execution checks that the return type of the answer is either a float or an int and verifies that there are no execution errors.
@@ 3.5 Expert-level Performance Evaluation
To provide a rough but informative estimate of the performance of domain-experts on each of DocMath-Eval sets, we invite two professionals with Chartered Financial Analyst licenses for evaluation. Regarding human expert performance on DM and DM, we report the same results as those in the original papers, with accuracy of 91% and 87%, respectively. For DM and DM, We randomly sample 25 examples from each set, asking the expert evaluators to answer the questions individually within a four-hour period. They achieve accuracy of 88% and 80% on DM (average 84%); and accuracy of 72% and 80% on DM (average 76%).
@@ 3.6 Dataset Release
Table 1 describes the basic statistics of four developed evaluation sets. DocMath-Eval contains a total of 4,000 questions with high-quality annotations, featuring varying difficulty levels in numerical reasoning and document understanding. We randomly divided the dataset into two splits: testmini and test. The testmini set contains 800 examples and is intended for model development validation. The test set comprises the remaining 3,200 examples and is designed for standard evaluation. To prevent data contamination Deng et al. (2024), the features relevant to the ground truth for the test set are not publicly released. Instead, we developed and maintain an online evaluation platform, allowing researchers to evaluate models and participate in a leaderboard.
@@ 4 Experiment Setup
4.1 Large Language Models
Our goal is to investigate the capabilities of current state-of-the-art LLMs on DocMath-Eval to better understand their strengths and limitations. To this end, we evaluate a wide range of models:
General: GPT-3.5&4 OpenAI (2022, 2023, 2024), Gemini-1.5 Gemini (2024), Claude-3&3.5 Anthropic (2024), Llama-2&3&3.1 Touvron et al. (2023); AI@Meta (2024), Mistral Jiang et al. (2023), Phi-3 Abdin et al. (2024), Gemma-1&2 Team et al. (2024), WizardLM-2 Xu et al. (2023), Yi-1.5 01.AI (2023), Qwen-2 Bai et al. (2023), Command R+ Cohere (2024b), Aya Cohere (2024a), and GLM-4 GLM et al. (2024).
We select the most recent checkpoint available as of August 1, 2024. The details of each evaluated model, including the exact model version we used, are presented in Appendix A in the Appendix.
@@ 4.2 Prompting Methods
Following recent LLM reasoning benchmark works Lu et al. (2024); Chen et al. (2023b), we evaluate two established prompting methods:
@@
Different from CoT, the PoT method Chen et al. (2023a) disentangles computation from the reasoning process by prompting the LLMs to generate a structured program to represent the reasoning process. The final answer is then derived by executing the generated program with an external calculator. Figure 3 in Appendix presents the PoT prompt used in our experiment.
@@
DeepSeek-Coder-V2 | 236B | Code | 85.0 | 79.0 | 78.0 | 66.5 | 56.0 | 54.0 | 41.0 | 37.7 | 63.1 | 57.3 | ||||
@@ 4.3 Implementation Details
The experiments for open-sourced LLMs were conducted using the vLLM framework Kwon et al. (2023). For all the experiments, we set temperature as 1.0 and maximum output length as 512. Given the extensive context length of input document, the main evaluation of DocMath-Eval is conducted under a zero-shot setting, aiming to assess LLMs’ capabilities to generate accurate answers without few-shot demonstrations or additional training.
@@
Building on previous work that evaluated LLMs on table-relevant tasks Chen (2023); Zhao et al. (2023b, c), we present our method for processing tabular data in documents. Specifically, we separate headers or cells in different columns using a vertical bar (|), and rows using a newline. This approach allows for the direct feeding of flattened table input into LLMs. In our preliminary study, we discovered that most LLMs can effectively understand such table representations. Nevertheless, we believe future research could explore more effective methods for encoding tabular data.
@@
For the DM subset, the input document length is extremely long and exceeds the context length limit of evaluated LLMs. Therefore, in our main experiments with DM, we evaluate models using the retrieval-augmented generation (RAG) setting. In this setting, external retrievers are employed to extract the top- most relevant textual and tabular evidence from the source document. We maintain the original relative order of the evidence and input it into the LLMs to answer the given question. We experiment with commonly-used (1) sparse retriever, i.e., BM25 Robertson et al. (1995), and (2) three dense retrievers, including OpenAI Embedding 3 small & large versions Neelakantan et al. (2022) and Contriever Izacard et al. (2022).
@@
For LLM with CoT prompting, we adopt the answer extraction pipeline from Chen et al. (2023b) to identify the final answer from the model’s output. For LLM with PoT prompting, we first develop a heuristic method to extract the generated python solution from the model’s output. If this python solution is executable, we execute it to obtain the final answer.
@@ 5.1 Main Results
Table 2 and Table 9 in the Appendix present the LLM performance using CoT and PoT prompting methods on the DocMath-Eval testmini and test sets, respectively. While the current best-performing LLM, GPT-4o, achieves performance comparable to human experts in simple problem settings (i.e., DM and DM), we find significant performance gaps in more challenging settings. Specifically, GPT-4o achieves an accuracy of 41.0% on DM with PoT, which is far behind the human expert performance of 76.0%. This underscores the need for ongoing LLM development, particularly in complex problem-solving over long and specialized documents.
Most open-source LLMs still lag behind the proprietary LLMs. However, the two DeepSeek-V2 models achieve performance levels close to the best-performing proprietary models. This indicates the potential for open-source LLMs to close the performance gap with proprietary models in the near future, given continued innovation and community collaboration. Additionally, the proprietary LLMs and code-specific models typically achieve comparable or better performance when using PoT prompting compared to CoT prompting. For math-specific LLMs, InternLM2-Math-Plus surpasses its backbone models in CoT performance, improving from 9.9% to 13.0% on average accuracy. This demonstrates the effectiveness of instruction-tuning in enhancing math reasoning.
@@ 5.2 Analysis on DM Set
In addition to using RAG for analyzing long specialized documents, recent advancements have extended the input length of LLMs to handle lengthy documents Su et al. (2023). We compare models with a context length limit of over 100K under both the RAG (as used in the main results) and Long-Context settings, where the entire document is input. As illustrated in Table 4, the evaluated models generally achieve close performance under RAG and long-context settings. This indicates that models with extended context lengths can effectively process lengthy inputs without a significant drop in performance compared to the RAG setting.
Error Analysis
To better understand the strengths and weaknesses of LLMs, we conduct an extensive error analysis. This analysis focuses on 100 randomly selected examples from the DM testmini set where GPT-3.5-turbo failed. We identify four common types of errors in current LLMs: inaccurate evidence retrieval, calculation errors, table misunderstandings, and exceeding context length. A detailed explanation for each type is provided in Appendix A in Appendix.
6 Conclusion
Limitations
In this work, we propose DocMath-Eval and conduct comprehensive analysis of different LLMs’ capabilities in solving knowledge-intensive math reasoning problems in finance domains. However, there are still some limitations: First, our method for extracting final answer from model output is still not perfect. In some cases, this methods fails to locate the answer, leading to the reported accuracy being an approximate lower bound. Moreover, due to computational resource constraints, we do not tune LLMs on a large-scale finance-domain data ourselves. However, we believe that training on finance data can help improve LLMs’ capabilities in solving problems in DocMath-Eval Wu et al. (2023b); Luukkonen et al. (2023); Xie et al. (2023).
Acknowledgement
We are grateful for the compute support provided by Microsoft Research’s Accelerate Foundation Models Research (AFMR) program. We would also like to thank the anonymous reviewers and area chairs for constructive discussions and feedback.
Appendix A Appendix
Annotation Quality | %S 4 |
Question Fluency | 97.4 |
Question Correctness | 96.0 |
[…]
Evidence Completeness | 91.3 |
[…]
Python Solution Correctness | 97.6 |
Variable Value Correctness | 98.5 |
Python Solution Conciseness | 89.1 |
Variable Name Meaningfulness | 95.4 |
Comment Comprehensiveness | 87.4 |
@@ Appendix A Appendix
Exceeding Context Length
(8 / 100) |
The input document exceeds the context length limit. |

### v3

@@ Abstract
Recent LLMs have demonstrated remarkable performance in solving exam-like math word problems. However, the degree to which these numerical reasoning skills are effective in real-world scenarios, particularly in expert domains, is still largely unexplored. This paper introduces DocMath-Eval, a comprehensive benchmark specifically designed to evaluate the numerical reasoning capabilities of LLMs in the context of understanding and analyzing specialized documents containing both text and tables. We conduct an extensive evaluation of 48 LLMs using Chain-of-Thought and Program-of-Thought prompting techniques, aiming to comprehensively assess the capabilities and limitations of existing LLMs in DocMath-Eval. We found that even the current best-performing system (i.e., GPT-4o) still significantly lags behind human experts in solving complex numerical reasoning problems grounded in long contexts. We believe that DocMath-Eval can serve as a valuable benchmark for evaluating LLMs’ capabilities in solving challenging numerical reasoning problems within expert domains.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have attracted significant attention due to their capabilities in solving a broad range of tasks OpenAI (2023); AI@Meta (2024), including math word problems (MWPs) commonly found in academic exams Wang et al. (2017); Miao et al. (2020); Amini et al. (2019); Cobbe et al. (2021); Hendrycks et al. (2021); Cobbe et al. (2021); Lu et al. (2023); Chen et al. (2023b). These MWPs vary from basic arithmetic to advanced algebra, showcasing LLMs’ proficiency in numerical reasoning — a crucial skill for interpreting and manipulating numerical data across various contexts. Despite this progress, there is still a significant gap in understanding the practicality of LLMs’ numerical reasoning in real-world scenarios, particularly in specialized fields such as finance, medicine, and science. As illustrated in Figure 1, these expert domains necessitate LLMs to interpret complex, domain-specific documents, applying numerical reasoning to complex problem-solving Chen et al. (2021); Zhu et al. (2021); Zhao et al. (2022); Li et al. (2022b). Recognizing this gap, our research focuses on the finance domain Li et al. (2022a); Wu et al. (2023a); Yang et al. (2023); Callanan et al. (2023); Xie et al. (2024). The finance industry often deals with lengthy and data-intensive documents that demand advanced numerical reasoning skills for accurate analysis and decision-making.
@@
We conduct an extensive evaluation on DocMath-Eval, covering a total of 48 proprietary and open-source LLMs from 17 organizations. Two prompting methods, Chain-of-Thought (CoT) Wei et al. (2022) and Program-of-Thought (PoT) Chen et al. (2023a), are applied for result analysis. Our experimental results indicate that while the existing best-performing LLM on average (i.e., GPT-4o) can achieve high performance in simple settings (e.g., DM), it still falls short of human experts in more challenging ones, i.e.,, DM. Moreover, Claude-3.5-Sonnet outperforms other LLMs, achieving an accuracy of 40.0% on the DM set when applying CoT prompting. However, it still lags far behind human expert performance, which stands at 76%. This significant gap between LLMs and human experts underscores the challenges presented by DocMath-Eval. It underscores the importance of advancing LLMs’ numerical reasoning and document understanding abilities to effectively apply them in the real-world specialized domains.
@@ 2 Related Work
Numerical reasoning over documents requires models to have a deep understanding of context and the ability to derive answers through numerical reasoning Dua et al. (2019). Applying these models in the finance domain Xie et al. (2023); Wu et al. (2023a); Yang et al. (2023) presents additional challenges in terms of interpreting hybrid data Zhu et al. (2021) and utilizing domain-specific expertise Chen et al. (2021); Zhao et al. (2024). Numerous datasets focusing on numerical reasoning over specialized documents have been proposed recently. Two notable benchmarks are TAT-QA Zhu et al. (2021) and FinQA Chen et al. (2021), which represent pioneering efforts in studying numerical reasoning in finance, particularly requiring the fusion of tabular and textual content. Building upon TAT-QA, a more challenging dataset named TAT-HQA Li et al. (2022b) was developed, focusing on counterfactual questions in relation to the provided context. Additionally, MultiHiertt Zhao et al. (2022) focuses on numerical reasoning over longer financial documents containing multiple tables. However, as illustrated in Table 1, these four datasets focus on less challenging scenarios, where either simple numerical reasoning (e.g., calculating the increasing rate or average value) is sufficient, or the input context is short. Furthermore, there is a lack of a standardized benchmark for systematically evaluating models’ performance across varying difficulty levels in terms of numerical reasoning and document understanding.
@@ 3.2 Solution Format Standardization
We observe that existing finance QA datasets feature solutions in various formats. Specifically, TAT-QA Zhu et al. (2021) and TAT-HQA Li et al. (2022b) utilize text, while MultiHiertt Zhao et al. (2022) employs mathematical expressions, such as 100/3, and FinQA Chen et al. (2021) uses math programs, such as divide(100,3), for solution annotations. This diversity in annotation formats hinders the development of a unified evaluation framework to assess LLM performance across different benchmarks. Additionally, text-based solutions often fall short in precision and clarity, making them less suitable for computational problem-solving; and the solutions presented as mathematical equations or programs can be less descriptive, with the intended semantic meaning of the equations sometimes being unclear.
To overcome the aforementioned limitations, in DocMath-Eval, we represent solutions using Python programs Zhao et al. (2024). Such a unified Python program format supports a standardized and effective evaluation framework for LLM assessment. Specifically, annotators are instructed to initially define variables at the start of the Python function, beginning with “def solution():”. These variables should align with the primary elements or quantities referenced in the question or relevant content in the documents. They then write a Python program that methodically address the problem, solving it step by step. Additionally, annotators receive a bonus for writing detailed comments, thereby enhancing the code’s readability and understandability. To verify the correctness and performance of the solutions, our annotation interface automatically runs the Python function. This process checks that the output is either a float or int and ensures that the execution finishes without any errors.
@@ 3.5 Expert-level Performance Evaluation
To give a general yet insightful estimate of the performance on each of the DocMath-Eval sets, we enlisted two professionals who hold Chartered Financial Analyst licenses to conduct the evaluation. Regarding human expert performance on DM and DM, we report the same results as those in the original papers, with accuracy of 91% and 87%, respectively. For DM and DM, We randomly sample 25 examples from each set, asking the expert evaluators to answer the questions individually within a four-hour period. They achieve accuracy of 88% and 80% on DM (average 84%); and accuracy of 72% and 80% on DM (average 76%).
@@ 3.6 Dataset Release
Table 1 presents the data statistics of four developed evaluation sets. DocMath-Eval contains a total of 4,000 questions with high-quality annotations, featuring varying difficulty levels in numerical reasoning and document understanding. We randomly partitioned the dataset into two subsets: testmini and test. The testmini subset includes 800 examples and is intended for model development and validation. The test subset consists of the remaining 3,200 examples, which are reserved for standard evaluation. To avoid data contamination Deng et al. (2024), the features directly related to the ground truth for the test set are kept private. Instead, we have developed and manage an online evaluation platform, where researchers can assess models and participate in a leaderboard.
@@ 4 Experiment Setup
4.1 Evaluated Large Language Models
Our goal is to investigate the capabilities of current state-of-the-art LLMs on DocMath-Eval to better understand their strengths and limitations. To this end, we evaluate a wide range of models, including 32 general-purpose LLMs, 4 math-specific LLMs, 6 code-based LLMs, and 7 mixture of experts (MoE) models. The specific details of each evaluated LLM, including the exact version used, can be found in Appendix A in the Appendix.
@@ 4.2 Prompting Methods
Following recent works on LLM reasoning benchmarks Lu et al. (2024); Chen et al. (2023b), we evaluate two commonly used prompting methods for math reasoning:
@@
The PoT method Chen et al. (2023a) separates computation from the reasoning process by instructing the LLMs to produce a structured program that encapsulates the reasoning steps. The final answer is obtained by executing the generated program. Figure 3 in Appendix presents the PoT prompt we used.
@@
| DeepSeek-Coder-V2 | 236B | Code, MoE | 85.0 | 79.0 | 78.0 | 66.5 | 56.0 | 54.0 | 41.0 | 37.7 | 63.1 | 57.3 | ||||
@@ 4.3 Implementation Details
The experiments involving open-sourced LLMs were conducted using the vLLM framework Kwon et al. (2023). In all the experiments, we used a temperature setting of 1.0 and maximum output length of 512. Given the extensive context length of input document, the main evaluation of DocMath-Eval is conducted under a zero-shot setting, aiming to assess LLMs’ capabilities to generate accurate answers without few-shot demonstrations or additional training.
@@
Building on previous work that evaluated LLMs on table-relevant tasks Chen (2023); Zhao et al. (2023a, b), we present our method for processing tabular data in documents. Specifically, we separate headers or cells in different columns using a vertical bar (|), and rows using a newline. This approach allows for the direct feeding of flattened table input into LLMs. In our preliminary study, we found that most LLMs can comprehend these table formats well. Nevertheless, we believe that future research could explore more effective methods for encoding tabular data Fang et al. (2024).
@@
For the DM subset, the input document length is extremely long and exceeds the context length limit of evaluated LLMs. Therefore, in our main experiments with DM, we evaluate models using the retrieval-augmented generation (RAG) setting. In this setting, external retrievers are employed to extract the top- most relevant textual and tabular evidence from the source document. We maintain the original relative order of the evidence and input it into the LLMs to answer the given question. We experiment with commonly-used sparse retriever, i.e., BM25 Robertson et al. (1995), and three dense retrievers, including OpenAI Embedding 3 small & large versions Neelakantan et al. (2022) and Contriever Izacard et al. (2022).
@@
For LLMs using CoT prompting, we adopt the answer extraction process developed by Chen et al. (2023b) to extract the final answer from the model’s output. For LLMs employing PoT prompting, we first develop a heuristic method to extract the generated python solution from the model response. We then execute it to obtain the final answer.
@@ 5 Results and Analysis
We next discuss our main findings from the experiments and our analysis of the DM subset.
@@ 5.1 Main Results
Table 2 and Table 9 in the Appendix present the LLM performance on the DocMath-Eval testmini and test sets, respectively.
While the current best-performing LLM, GPT-4o, achieves performance comparable to human experts in simple problem settings (i.e., DM and DM), we find significant performance gaps in more challenging settings. Specifically, GPT-4o achieves an accuracy of 41.0% on DM with PoT, which is far behind the human expert performance of 76.0%. This underscores the need for ongoing LLM development, particularly in complex problem-solving over long and specialized documents. Most open-source LLMs still lag behind the proprietary LLMs. However, the two DeepSeek-V2-* models come close to matching the performance of the leading proprietary models. The DeepSeek-V2 even outperforms GPT-4o on the DM subset. This suggests that open-source LLMs have the potential to bridge the performance gap with the leading proprietary models in the near future.
The code-specific and proprietary LLMs generally perform as well as or better with PoT prompting compared to CoT prompting. This is likely because LLMs are prone to making errors during complex mathematical computations, as revealed in concurrent work Zhao et al. (2024). Additionally, for math-specific LLMs, InternLM2-Math-Plus outperforms its base model in CoT performance, with average accuracy rising from 9.9% to 13.0%. This highlights the impact of instruction-tuning in improving math reasoning abilities.
@@ 5.2 Analysis on DM Set
In addition to using RAG for analyzing long specialized documents, recent advancements have extended the input length of LLMs to handle lengthy documents Su et al. (2023). We compare models with a context length limit of over 100K under both the RAG (as used in the main results) and Long-Context settings, where the entire document is input. As illustrated in subsection 5.2, the evaluated models generally achieve close performance under RAG and long-context settings. This indicates that models with extended context lengths can effectively process lengthy inputs without a significant drop in performance compared to the RAG setting.
@@
Error Analysis
To better understand the strengths and weaknesses of LLMs, we conduct an extensive error analysis. This analysis focuses on 100 randomly selected examples from the DM testmini set where GPT-3.5-turbo failed. We identify four common types of errors in current LLMs: inaccurate evidence retrieval, calculation errors, table misunderstandings, and exceeding context length. A detailed explanation for each type is provided in subsection 5.2.
6 Conclusion
Limitations
There are some limitations in our study that we believe can be addressed in future work. First, our approach to extracting the final answer from the model’s output is not yet flawless. In certain instances, this method fails to accurately identify the answer, causing the reported accuracy to be an approximate lower limit. Additionally, we suggest that future research could investigate training large language models (LLMs) on finance-specific data to improve their performance on the DocMath-Eval benchmark Wu et al. (2023b); Luukkonen et al. (2023); Xie et al. (2023).
Acknowledgement
We are grateful for the compute support provided by Microsoft Research’s Accelerate Foundation Models Research (AFMR) program. We extend our gratitude to the anonymous reviewers and area chairs for their valuable discussions and feedback.
Appendix A Appendix
| Annotation Quality | %S 4 |
| Question Fluency | 97.4 |
| Question Correctness | 96.0 |
[…]
| Evidence Completeness | 91.3 |
[…]
| Python Solution Correctness | 97.6 |
| Variable Value Correctness | 98.5 |
| Python Solution Conciseness | 89.1 |
| Variable Name Meaningfulness | 95.4 |